a theater got 4275$ for 675 people, for adults it's 7$ and for children 5$ how many people (adults and kids) attend that night?
Answer by Boreal(15235)

x adults
675-x children
7x+5(675-x)=4275
7x+3375-5x=4275
2x=900
x=450 adults (3150)
675-x=225 children (1125)